OpenStudy (anonymous):

Find the sum of 4x^3 + 2x^2 - 4x + 3 and 7x^3 -4x^2 + 7x + 8

OpenStudy (raden):

just combine the similar terms, like 4x^3 + 7x^3 = 11x^3 and so on ...

OpenStudy (anonymous):

??

OpenStudy (raden):

can u find the other which has the similar terms and add them

OpenStudy (raden):

2x^2 + (-4x^2) = 2x^2 - 4x^2 = ...

OpenStudy (raden):

-4x + 7x = ...

OpenStudy (raden):

3+8 = ...

OpenStudy (raden):

finally, joining all sum above